An astronomer is interested in measuring in light years, the distance from his observatory to a distant star. If the astronomer believes that the values of the measurements are independent and identically distributed random variables with common mean d (the actual distance) and a common variance of 9 (light years), how many measurements must he make to be 99% sure that his estimated distance is accurate to within + 1.0 light year? (Note: for 95% confidence, Phi(Z) >-0.975, which means that Z-1.96) (Note: for 99% confidence, phi(Z/>-0.995, which means that Z-2.575) o n 2 [106.09] 768.32 n137.26 n>「245.861 n-(59.68) 61.47 n96.04 34.57
Answer #1

here we have,

for 99% of CI, Z = 2.575

standard deviation, SD = sqrt(variance) = sqrt(9) = 3

margin of error, E = 1

required sample size, n = (Z*SD/E)2 = (2.575*3/1)2 = 59.68

correct option is n>=59.68